# Absence of gapless Majorana edge modes in few-leg bosonic flux ladders

### v1

@@ Abstract
We perform DMRG simulations [30, 31, 32, 33] to find an accurate matrix product state (MPS) representation of the ground state [34]. From the MPS, we extract the bipartite entanglement entropy $S(\ell)$ as a function of the cut position $\ell$ along the long $x$-direction of the system. The entanglement entropy $S(\ell)$ is related to the central charge $c$ via Calabrese and Cardy’s CFT prediction [35],
@@
Figure 3: (a) Central charge $c$ as extracted from the bipartite entanglement entropy using a scaling collapse for $L_{y}\!=\!4$ and $N_{\rm max}\!=\!4$ without local three-body repulsion. We find the same phases as for the case $N_{\rm max}\!=\!2$ (see Fig. 1). (b) This is also evident in the isotropic limit, $t_{y}\!=\!t_{x}$, where the central charges for simulations with ($N_{\rm max}\!=\!2$, red circles) and without ($N_{\rm max}\!=\!4$, blue diamonds) hard-core three-body repulsion are in excellent agreement. In both panels the gray region indicates very weak interaction strengths $U/t_{x}<0.5$ where for $N_{\rm max}\!=\!4$ the central charge could not be determined with the necessary accuracy.
@@
In such small systems the edge will exhibit strong finite-size effects, so that we do not necessarily expect to find gapless edge modes [23]. Instead, to analyze the nature of the many-body ground state, we have calculated its particle entanglement spectrum (PES) for a particle partition with $N_{A}\!=\!N_{B}\!=\!2$. For a fractional quantum Hall ground state, the PES [21, 22] is gapped, and the number of states below the gap is equal to the degeneracy of quasihole states for a system with $N_{A}$ bosons in the same geometry, and is therefore a universal indicator of the nature of the ground state.
@@
Figure 4: Stability regime of the bosonic Pfaffian state in short ladders, as extracted through the particle entanglement spectrum of 4-boson ground states obtained through exact diagonalization. The yellow region indicates a finite entanglement gap with both particle cuts $N_{A}\!=\!1$ and $N_{A}\!=\!2$, with a counting consistent with the 4 particle Pfaffian (respectively 3 and 6 states below the entanglement gap). The ladder lengths (respectively $L_{x}\!=\!12,10,9,8$) are adjusted to the number of legs $L_{y}$ to have a sufficient number of ladder sites while keeping the computational complexity manageable.

### v2

@@ Abstract
We perform DMRG simulations [30, 31, 32, 33] to find an accurate matrix product state (MPS) representation of the ground state. We exploit the $\mathrm{U}(1)$-symmetry associated with particle number conservation and consider bond dimensions up to $\chi=6000$ [34]. From the MPS, we extract the bipartite entanglement entropy $S(\ell)$ as a function of the cut position $\ell$ along the long $x$-direction of the system. The entanglement entropy $S(\ell)$ is related to the central charge $c$ via Calabrese and Cardy’s CFT prediction [35],
@@
Figure 3: (a) Central charge $c$ as extracted from the bipartite entanglement entropy using a scaling collapse for $L_{y}\!=\!4$ and $N_{\rm max}\!=\!4$ without local three-body repulsion. We find the same phases as for the case $N_{\rm max}\!=\!2$ (see Fig. 1). (b) This is also evident in the isotropic limit, $t_{y}\!=\!t_{x}$, where the central charges for simulations with ($N_{\rm max}\!=\!2$, red circles) and without ($N_{\rm max}\!=\!4$, blue diamonds) hard-core three-body repulsion are in excellent agreement. In both panels the gray region indicates very weak interaction strengths $U/t_{x}<0.5$ where for $N_{\rm max}\!=\!4$ the central charge could not be determined with the necessary accuracy. Ladder lengths used for the scaling collapse: $L_{x}\!=\!25,33,49,61$.
@@
In such small systems the edge will exhibit strong finite-size effects, so that we do not necessarily expect to find gapless edge modes [23]. Instead, to analyze the nature of the many-body ground state, we have calculated its particle entanglement spectrum (PES) for a particle partition with $N_{A}\!=\!N_{B}\!=\!2$. For a fractional quantum Hall ground state, the PES [21, 22] is gapped, and the number of states below the gap is equal to the degeneracy of quasihole states for a system with $N_{A}$ bosons in the same geometry, and is therefore a universal indicator of the nature of the ground state. Being an integer by construction, the number of states below the gap is less prone to finite-size effects than other quantities obtained numerically, like the many-body Chern number [46] or the central charge.
@@
Figure 4: Stability regime of the bosonic Pfaffian state in short ladders, as extracted through the particle entanglement spectrum of 4-boson ground states obtained through exact diagonalization. The colormap indicates the entanglement gap above the lowest $6$ states for a particle cut with $N_{A}\!=\!2$. The red black lines indicates a finite entanglement gap with both particle cuts $N_{A}\!=\!1$ and $N_{A}\!=\!2$, with a counting consistent with the 4 particle Pfaffian (respectively 3 and 6 states below the entanglement gap). The ladder lengths (respectively $L_{x}\!=\!12,10,9,8$) are adjusted to the number of legs $L_{y}$ to have a sufficient number of ladder sites while keeping the computational complexity manageable.